doing equations with 3 unknowns having a hard time with this any help would be appreciated
3x+3y=3
-3x-2z=-11
y+z=-1

Answer by lwsshak3(11628)

(eq1) 3x+3y=3
(eq2)-3x-2z=-11
(eq3)y+z=-1
***
3x+3y=3 (eq1)
-3x-2z=-11(eq2)
add:
3y-2z=-8 (eq4)
..
y+z=-1 (eq3)
2y+2z=-2 (mult. eq by 2)
3y-2z=-8 (eq4)
add:
5y=-10
y=-2
z=-1-y=-1+2=1
3x=3-3y=3-3(-2)=9
x=3
Check:
***
3x+3y=3 (eq1)
-3x-2z=-11(eq2)
add:
3y-2z=-8 (eq4)
..
y+z=-1 (eq3)
2y+2z=-2 (mult. eq by 2)
3y-2z=-8 (eq4)
add:
5y=-10
y=-2
z=-1-y=-1+2=1
3x=3-3y=3-3(-2)=9
x=3
y=-2
z=1
Check:
3x+3y=3*3+3*-2=9-6=3
-3x-2z=-3*3-2*1=-9-2=-11
y+z=-2+1=-1
check ok: